### v1

@@ 3.3 ViT-based methods
Figure 7: General structure of ViT, where an asterisk (*) indicates an additional learnable class embedding. The multilayer perceptron (MLP) is responsible for feature transformation, global context fusion, and classification or regression.
@@ 4 LLM-based methods
Figure 9: A taxonomy categorizing LLMs into three distinct groups.
@@ 4.1 Encoder-only-based methods
The system log produced by a computer system comprises extensive data gathered simultaneously, serving as the foundational data for identifying errors, intrusions, and abnormal behaviors. Detecting anomalies in system logs aims to swiftly identify irregularities with minimal human intervention, a critical challenge in the industry. An approach previous suggested typically involved converting log data into a standardized template using a parser before anomaly detection algorithms could be applied. For instance, LogBERT [84], a BERT-based anomaly detection framework, refines log sequences using a drain parser. It trains solely on normal log data through two tasks: masked log key prediction and volume of hypersphere minimization. The first task trains normal log patterns using masked language modeling, while the second task identifies the smallest sphere containing normal logs. During inference, top predicted log keys form a candidate set from a randomly masked normal log sequence. An observed log key not belonging to this set indicates an anomaly. However, LogBERT’s random log key selection during masking limits its ability to fully consider the log sequence. Defining templates for specific events in advance could lead to loss of information within the log keys. In the study conducted by Lee et al. [80], they introduced LAnoBERT, a parser-free method for system log anomaly detection that leverages the BERT model, demonstrating strong NLP capabilities. LAnoBERT learns the model through masked language modeling, a BERT-based pre-training method, and employs unsupervised learning-based anomaly detection using the masked language modeling loss function per log key during testing. Experiments demonstrated that LAnoBERT not only outperforms unsupervised learning-based benchmark models in anomaly detection but also achieves comparable performance to supervised learning-based benchmark models. Figure 10 depicts the proposed Transformers and LLM-based intrusion and anomaly detection, a modified version of BERT or Attention Transformers, serving as the foundation of an IDS for such models. The figure highlights the essential components of Transformers and LLMs, providing a conceptual representation focused on the core components. An LLM model, such as BERT, can be substituted with one of the Transformer or LLM models indicated in Figure 10 and paired with a decoder from Figure 9. This configuration yields many new encoder-decoder LLMs, developed in accordance with the BERT2GPT principle, specifically for IDS applications. Additionally, merging model blocks from the same figure could produce a specialized encoder dedicated to IDS.
@@ 4.2 Decoder-only-based methods
Web fuzzing is a security testing technique that generates and sends random or mutated inputs to a web application to discover vulnerabilities. It uses coverage feedback from the application to refine its inputs, helping identify potential security flaws. The work proposed in [86] introduced GPTFuzzer, which employs an encoder-decoder architecture to produce effective payloads for web application firewalls, specifically targeting SQL injection, XSS, and remote code execution (RCE) attacks by generating fuzz test cases. This model undergoes reinforcement learning (RL) [87], TL and applies a Kullback-Leibler divergence (KLD) [3] penalty to efficiently generate attack payloads and avoid local optimum issues. Similarly, [88] utilized an encoder-decoder architecture model to create SQL injection detection test cases for web applications, translating user inputs into new test cases.
@@
[37] CAN-BERT Identify cyber threats on the CAN network. Requires substantial computational demands and memory usage. LSTM-AE In-vehicle F1= 81-99 F1 $\approx$ +10
@@ 5 Environments and applications for Transformers- and LLM-based IDS
This section explores the diverse applications of Transformer- and LLM-based IDS. Spanning across computer networks, IoT, critical infrastructure, cloud and SDN, as well as AV and unmanned unmanned aerial vehicles (UAV), these environments showcase the adaptive capabilities of modern IDS technologies. Highlighting their relevance in safeguarding diverse systems, this section delves into their efficacy and adaptability in various complex and dynamic environments. Besides, Table 6 outline the effectiveness and constraints of particular environments and applications in current SOTA Transformer and LLM-based IDS implementations.
@@ 5.2 Internet-of-things
NLP techniques, encompassing Transformers and LLM, have already been employed as a means to build robust and efficient IDS methods in IoT environments. TransIDS [114], a Transformer-based method, adaptively focuses on important features for IoT intrusion detection. Moving on, the study [115] introduces a Transformer neural network-based IDS (TNN-IDS) for message queuing telemetry transport (MQTT)-enabled IoT networks. TNN-IDS addresses limitations from imbalanced training data by leveraging parallel processing and MHA layers in the Transformer, enhancing the detection and learning of malicious activities. In smart home IoT scenarios with fewer devices, both network traffic-based and telemetry data-based NIDS can work independently. This motivated the researchers to propose a novel ML-based NIDS that combines these approaches in [116]. They introduce a Transformer-based IoT NIDS, which includes self-Attention, to learn attack behaviors from diverse data in heterogeneous IoT environments. Other researchers have explored similar techniques for IIoT, as documented in references [53, 54, 117, 118], akin to previous studies. However, [118] utilizes micro-controller unit (MCU) temperature fingerprints for IIoT intrusion detection. It records temperature sequences, analyzes their relationship with node complexity, computes residuals, and employs a self-encoder model for security assessment. Addressing the escalating challenge of network security, Yan et al. [117] propose a multi-Transformer fusion intrusion detection model tailored for real-world IIoT environments.
@@
[…]
[119] 2022 MHA Network Attacks present in CICIDS2017 and CIC-DDoS2019 datasets F1= 99.17, 98.48 "Real-time detection has not been verified." NA
@@
[128] 2023 MHA AV Malfunction attack Acc= 99.75 Generalizability unverified due to testing on a single dataset NO
Abbreviations: LLM or Transformer used (LoTU), Deep learning model (DLM); Best performance (BP); Availability of data and materials (ADM), Upon request (UR), not available (NA).
@@ 5.5 Autonomous vehicles
AV face increasing vulnerability to intrusion threats. Addressing these challenges, leveraging LLMs and Transformers technologies becomes crucial. These advanced models enhance detection and response capabilities, safeguarding these systemss from evolving cybersecurity risks. Specifically, AVs infrastructures heavily rely on sensor and electronic component signals, facilitated by wireless technologies that enhance communication but also increase vulnerability to malicious disruptions. The paper [128] proposes a Transformer neural network-based intrusion detection system (CAN-Former IDS) to predict anomalies in CAN protocol communications, addressing both the sequence of IDs and message payload values. Advantages include fully self-supervised training and token interaction without hand-crafted features, evaluated using the survival analysis dataset with CAN communication from three vehicles. Similarly, Lai study in [129], introduces a federated learning-based edge computing (FL-EC) architecture to enhance privacy and security in vehicle-to-everything (V2X) communications within the Internet over vehicles. By enabling collaborative learning among edge devices without centralizing sensitive data, the FL-EC architecture addresses the limitations of traditional vehicles-to-cloud (V2C) systems. The study also presents the feature select Transformer (FSFormer), which employs a feature Attention mechanism to dynamically select significant features for IDS, improving the model’s ability to extract critical information. Experiments using the UNSW-NB15 and CSE-CIC-IDS2018 datasets showed that FSFormer achieved high accuracy and F1 scores, outperforming other baseline models and DL methods. This demonstrates FSFormer’s effectiveness in detecting intrusions and securing V2X communications.
@@ 7.3 Perspectives and future directions
Since LLMs represent the latest advancement in Transformer technology, future research and perspectives mostly focus exclusively on LLMs. Despite extensive research into LLMs within cybersecurity, their exploration and application are still in the early stages, offering considerable potential for growth [136]. The complexity of cybersecurity arises not only from the variety of attack methods but also from the intricate nature of network environments, combined with the necessity for a comprehensive application of diverse tools and strategies to ensure effective protection [137, 21]. Addressing these challenges requires AI systems with advanced capabilities in planning, reasoning, tool use, memory, and more. Building upon this discussion, the authors identify several critical perspectives that need to be addressed in LLM-based IDS, as follows:
@@
(g) Integration with dynamic security Ecosystems and adversarial resilience: Future LLMs for IDS need to increasingly integrate with dynamic security ecosystems, incorporating real-time threat intelligence from diverse sources such as threat feeds, vulnerability databases, and incident reports. This integration not only enriches model training but also enhances adversarial resilience by enabling LLMs to detect and mitigate sophisticated adversarial attacks. Additionally, collaborative defense systems where multiple specialized LLMs work synergistically will further strengthen IDS capabilities, improving overall threat detection accuracy and reducing false positives. This holistic approach ensures that LLM-based IDS are well-equipped to handle evolving cybersecurity challenges effectively.
(h) Cross-domain applications of Transformer and LLM-based IDS: The future direction of utilizing Transformer and LLMs in IDS presents promising cross-domain applications, notably in sectors like healthcare, finance, and other fields that encounter significant network security challenges. These industries are crucial due to their sensitivity and the severe consequences of data breaches. Integrating advanced IDS technologies based on Transformers or LLMs could vastly improve the detection and prevention of sophisticated cyber-attacks in these areas. For instance, in healthcare, leveraging these models can help in detecting anomalies in network traffic that could signify breaches of patient data. Similarly, in the financial sector, such systems can be trained to identify patterns indicative of fraud or data theft, providing an added layer of security against cyber threats. Moreover, the adaptability of Transformers and LLMs allows for customized solutions tailored to the specific security requirements of each field, enhancing overall system robustness. By extending the application of these advanced models beyond their typical use cases, researchers can create a more secure and reliable digital infrastructure across various critical domains.
@@
(i) Interpretability and scalability improvement: Integrating domain knowledge into Transformer and LLM-based IDS offers a promising direction for future research. By incorporating specific insights and contextual information from various fields, the interpretability of these models can be significantly improved. This approach allows for a deeper understanding of the patterns and anomalies detected, making the results more meaningful and actionable for cybersecurity experts. For instance, in the healthcare sector, integrating medical domain knowledge can help the model distinguish between normal and abnormal network activities related to medical devices and patient data flows. Similarly, in the financial sector, understanding the typical transaction patterns and identifying deviations can enhance the model’s ability to detect fraudulent activities.

### v2

@@ 3.4 ViT-based methods
Figure 7: General structure of ViT-based IDS, where an asterisk (*) indicates an additional learnable class embedding. The multilayer perceptron (MLP) is responsible for feature transformation, global context fusion, and classification or regression.
@@
De Rose et al. in [106] proposed a VINCENT method utilizes a ViT as the teacher model in its knowledge distillation setup to detect and classify cyber threats. It incorporates an imagery representation of cyber-data, where the ViT processes these images to generate attention-based features that are significant for classifying different types of threats. This methodology allows VINCENT to effectively recognize and categorize various cyber threats, achieving a notable performance with the best overall accuracy of 87.2%. However, the batch-based learning approach which does not account for the evolving nature of cyber threats, potentially limiting its effectiveness over time as new threat patterns emerge. Likewise, the DE-VIT scheme proposed in [107] leverages ViT capabilities for image classification techniques to tackle intrusion detection by transforming cyber threat data into a visual format. This method uses deformable attention mechanisms to focus selectively on parts of the data, rather than processing entire feature sets indiscriminately, thereby improving detection accuracy, especially in varied and complex datasets. The best result was achieved with a 99.5% accuracy on the CIC IDS2017 dataset. However, a notable limitation of DE-VIT is that its performance decreases rapidly in data categories with fewer samples, highlighting an issue with handling imbalanced datasets. Unlike previous approaches, Yang et al. [108] proposed the ConViTML approach, which integrates CNN and ViT to capture both local and global features in network traffic data. The CNN efficiently extracts low-dimensional, shallow features from individual packets, while the ViT captures the structural and sequential relationships between these packets. This fusion significantly enhances the model’s capability to classify encrypted malicious traffic with high precision. The best performance was achieved an accuracy of 99.75%. Even so, the model’s complexity and reliance on extensive training data may limit its scalability and real-time deployment efficiency.
@@ 4 LLM-based methods
Figure 9: A taxonomy categorizing LLMs into three distinct groups. The LLMs marked in bold are those already utilized for IDS and reviewed in this survey.
@@ 4.1 Encoder-only-based methods
The system log produced by a computer system comprises extensive data gathered simultaneously, serving as the foundational data for identifying errors, intrusions, and abnormal behaviors. Detecting anomalies in system logs aims to swiftly identify irregularities with minimal human intervention, a critical challenge in the industry. An approach previous suggested typically involved converting log data into a standardized template using a parser before anomaly detection algorithms could be applied. For instance, LogBERT [116], a BERT-based anomaly detection framework, refines log sequences using a drain parser. It trains solely on normal log data through two tasks: masked log key prediction and volume of hypersphere minimization. The first task trains normal log patterns using masked language modeling, while the second task identifies the smallest sphere containing normal logs. During inference, top predicted log keys form a candidate set from a randomly masked normal log sequence. An observed log key not belonging to this set indicates an anomaly. However, LogBERT’s random log key selection during masking limits its ability to fully consider the log sequence. Defining templates for specific events in advance could lead to loss of information within the log keys. In the study conducted by Lee et al. [112], they introduced LAnoBERT, a parser-free method for system log anomaly detection that leverages the BERT model, demonstrating strong NLP capabilities. LAnoBERT learns the model through masked language modeling, a BERT-based pre-training method, and employs unsupervised learning-based anomaly detection using the masked language modeling loss function per log key during testing. Experiments demonstrated that LAnoBERT not only outperforms unsupervised learning-based benchmark models in anomaly detection but also achieves comparable performance to supervised learning-based benchmark models. Moving forward, the paper in [117], proposes the LogFiT method, an advanced log anomaly detection model using a fine-tuned BERT-based language model. Unlike previous methods such as DeepLog and LogBERT, which rely on log templates for analysis, LogFiT operates directly on raw logs without requiring these templates, significantly improving anomaly detection accuracy through enhanced semantic analysis. The model leverages a self-supervised training regime, utilizing masked sentence prediction to deeply understand normal operational log data, making it highly effective in spotting deviations indicative of system threats. However, the method’s limitation includes high computational demands and a dependency on extensive pre-training.
Figure 10 depicts the proposed Transformers and LLM-based intrusion and anomaly detection, a modified version of BERT or Attention Transformers, serving as the foundation of an IDS for such models. The figure highlights the essential components of Transformers and LLMs, providing a conceptual representation focused on the core components. An LLM model, such as BERT, can be substituted with one of the Transformer or LLM models indicated in Figure 10 and paired with a decoder from Figure 9. This configuration yields many new encoder-decoder LLMs, developed in accordance with the BERT2GPT principle, specifically for IDS applications. Additionally, merging model blocks from the same figure could produce a specialized encoder dedicated to IDS.
@@ 4.2 Decoder-only-based methods
Web fuzzing is a security testing technique that generates and sends random or mutated inputs to a web application to discover vulnerabilities. It uses coverage feedback from the application to refine its inputs, helping identify potential security flaws. The work proposed in [84] introduced GPTFuzzer, which employs an encoder-decoder architecture to produce effective payloads for web application firewalls, specifically targeting SQL injection, XSS, and remote code execution (RCE) attacks by generating fuzz test cases. This model undergoes reinforcement learning (RL) [119, 120], TL and applies a Kullback-Leibler divergence (KLD) [3] penalty to efficiently generate attack payloads and avoid local optimum issues. Similarly, [69] utilized an encoder-decoder architecture model to create SQL injection detection test cases for web applications, translating user inputs into new test cases.
@@
The work recently proposed by Li et al. [127] introduced the DoLLM model, which enhances DDoS detection using LLM by structuring non-contextual network flows into sequential data that LLMs can process effectively. Specifically, the DoLLM method organizes flows into contextual sequences, converts them into tokens, and leverages the LLM’s semantic space to enhance detection accuracy, focusing on the potent Llama2-7B model for its robust processing capabilities. The approach significantly advances the detection of complex DDoS attacks, like carpet bombing, by effectively capturing and analyzing network flow data. The best result obtained by the DoLLM model is a high F1 score of 96.4% in zero-shot scenarios, indicating robust detection capabilities. However, the DoLLM method suffers from high computational overhead, reliance on external processing, and challenges in real-time detection due to latency.
@@
[37] CAN-BERT Identify cyber threats on the CAN network. Requires substantial computational demands and memory usage. LSTM-AE In-vehicle F1= 81–99 F1 $\approx$ +10 (Spark)
@@
[57] GPT-Neo Differentiating between normal and abnormal behaviors using an unsupervised learning method. However, the method reliance on predefined rules for log parsing, and computational costs. GPT-2 Web server log F1= 99.25 F1=+2.92
@@
[129] BigBird Integrating LLMs with system call analysis provides an enhanced malware detection capabilities. However, it lacks external processing independence and adaptation for detecting zero-day attacks with benign data. BERT, DistilBERT, GPT-2, Mistral, Longformer MalwSpecSys Acc= 86.67 Acc=+ 0.51
@@ 5 Environments and applications for Transformers- and LLM-based IDS
This section explores the diverse applications of Transformer- and LLM-based IDS. Spanning across computer networks, IoT, critical infrastructure, cloud and SDN, as well as AV and unmanned unmanned aerial vehicles (UAV), these environments showcase the adaptive capabilities of modern IDS technologies. Highlighting their relevance in safeguarding diverse systems, this section delves into their efficacy and adaptability in various complex and dynamic environments. Besides, Table LABEL:tab:Apps outline the effectiveness and constraints of particular environments and applications in current SOTA Transformer and LLM-based IDS implementations.
@@ 5.1 Computer network
Manocchio et al. [140] introduced the FlowTransformer framework, a transformative approach for implementing Transformer-based NIDS. This framework capitalized on the capabilities of Transformer models like GPT 2.0 and BERT to process complex patterns in network traffic, significantly enhancing the detection of sophisticated network threats. GPT 2.0, as a generative model, and BERT, as an encoder model, were evaluated for their efficacy in this domain, using various network datasets. The principle involved preprocessing network flow data into a format suitable for transformers, analyzing it through configurable transformer components, and efficiently classifying potential threats using a "Last Token" classification head. Notably, the model size could be reduced by over 50%, and both inference and training times were improved. The best performance achieved was an F1 score of 96.9. Yet, the method’s suffer from high dependency on data preprocessing and the challenge of applying appropriate Transformer configurations to network flow data.
@@ 5.2 Internet-of-things
NLP techniques, encompassing Transformers and LLM, have already been employed as a means to build robust and efficient IDS methods in IoT environments. TransIDS [141], a Transformer-based method, adaptively focuses on important features for IoT intrusion detection. Moving on, the study [74] introduces a Transformer neural network-based IDS (TNN-IDS) for message queuing telemetry transport (MQTT)-enabled IoT networks. TNN-IDS addresses limitations from imbalanced training data by leveraging parallel processing and MHA layers in the Transformer, enhancing the detection and learning of malicious activities. In smart home IoT scenarios with fewer devices, both network traffic-based and telemetry data-based NIDS can work independently. This motivated the researchers to propose a novel ML-based NIDS that combines these approaches in [62]. They introduce a Transformer-based IoT NIDS, which includes self-Attention, to learn attack behaviors from diverse data in heterogeneous IoT environments. Other researchers have explored similar techniques for IIoT, as documented in references [54, 55, 142, 67], akin to previous studies. However, [67] utilizes micro-controller unit (MCU) temperature fingerprints for IIoT intrusion detection. It records temperature sequences, analyzes their relationship with node complexity, computes residuals, and employs a self-encoder model for security assessment. Addressing the escalating challenge of network security, Yan et al. [142] propose a multi-Transformer fusion intrusion detection model tailored for real-world IIoT environments. Moving forward, Ferrag et al. [143] presented SecurityBERT, a novel BERT-based lightweight architecture tailored for cyber threat detection in IoT and IIoT networks. Utilizing a privacy-preserving encoding technique and the Byte-Pair Encoding Tokenizer, SecurityBERT efficiently processed network traffic data, including various IoT device metrics and network activity logs. The method leveraged BERT’s contextual understanding to enhance detection capabilities in IoT environments, making it highly suitable for deployment on resource-constrained devices. SecurityBERT achieved impressive performance metrics, setting new benchmarks for real-time traffic analysis in IoT devices. For intrusion prediction, Diaf et al. [144] introduced a novel network intrusion prediction framework designed for IoT networks, leveraging the strengths of GPT, BERT, and LSTM models. The principle of the proposed method involved using GPT to predict the next network packets based on current ones, BERT to evaluate the validity of these predictions, and LSTM to classify the packets as normal or malicious. The integration of these models enhanced the framework’s predictive capabilities, making it highly effective for IoT cybersecurity.
@@
[38] 2024 BERT IoV All attacks exists Car-Hacking, and in-Vehicle IDS datasets, and others. Acc= 99.97, 99.96 There is a need for more efficient and lightweight BERT models due to limited vehicular resources. NA
@@
[75] 2022 MHA Network Attacks present in CICIDS2017 and CIC-DDoS2019 datasets F1= 99.17,
98.48
"Real-time detection has not been verified." NA
@@
[82] 2023 MHA AV Malfunction attack Acc= 99.75 Generalizability unverified due to testing on a single dataset NA
[…]
[143] 2024 BERT IoT, IIoT All attacks present in Edge-IIoTset dataset Acc= 98.2 Generalizability is not verified due to training on one dataset. NA
[144] 2024 BERT
GPT-2
IoT All the attacks are present in the CICIoT2023 dataset Acc= 98 The model’s generalizability remains unverified because it was trained on a single dataset. NA
[147] 2024 MHA Satellite Attacks presents in UNSW-NB-15 and TON_IoT datasets Acc= 100 The method lacks a proper dataset that accurately mimics real satellite system data. NA
Abbreviations: LLM or Transformer used (LoTU), Inference time (IT); Model size (MS); Deep learning model (DLM); Best performance (BP); Availability of data and materials (ADM), Upon request (UR), not available (NA).
@@ 5.5 Autonomous vehicles
AV face increasing vulnerability to intrusion threats. Addressing these challenges, leveraging LLMs and Transformers technologies becomes crucial. These advanced models enhance detection and response capabilities, safeguarding these systemss from evolving cybersecurity risks. Specifically, AVs infrastructures heavily rely on sensor and electronic component signals, facilitated by wireless technologies that enhance communication but also increase vulnerability to malicious disruptions. The paper [82] proposes a Transformer neural network-based intrusion detection system (CAN-Former IDS) to predict anomalies in CAN protocol communications, addressing both the sequence of IDs and message payload values. Advantages include fully self-supervised training and token interaction without hand-crafted features, evaluated using the survival analysis dataset with CAN communication from three vehicles. Similarly, Lai study in [148], introduces a federated learning-based edge computing (FL-EC) architecture to enhance privacy and security in vehicle-to-everything (V2X) communications within the Internet over vehicles. By enabling collaborative learning among edge devices without centralizing sensitive data, the FL-EC architecture addresses the limitations of traditional vehicles-to-cloud (V2C) systems. The study also presents the feature select Transformer (FSFormer), which employs a feature Attention mechanism to dynamically select significant features for IDS, improving the model’s ability to extract critical information. Experiments using the UNSW-NB15 and CSE-CIC-IDS2018 datasets showed that FSFormer achieved high accuracy and F1 scores, outperforming other baseline models and DL methods. This demonstrates FSFormer’s effectiveness in detecting intrusions and securing V2X communications. The paper [38] presented IoV-BERT-IDS, a hybrid NIDS for the IoV in both in-vehicle and extra-vehicle networks, leveraging the BERT model. The proposed method involved a semantic extractor that transformed network traffic data, including CAN packets, into a format suitable for BERT processing. The model then underwent pre-training and fine-tuning to effectively detect intrusions. BERT played a critical role in capturing bidirectional contextual features, significantly improving the model’s generalization and accuracy in IoV environments.
@@ 6 Real-world case studies of Transformers and LLM-based IDSs
Deploying Transformer and LLM-based IDS in real-world scenarios presents a set of obstacles that can significantly affect their efficacy and feasibility. For instance, the complexity and computational demands of these models, as exemplified in the case studies using VMware setups and simulated attacks, indicate substantial hardware requirements. Such IDS implementations often necessitate robust servers and a substantial computer network and information technology infrastructure, which may not be viable for all organizations, especially smaller ones with limited budgets. Additionally, the real-time processing capabilities of these systems are crucial for effective network security but can be hampered by the hardware-intensive nature of Transformer and LLMs algorithms. For example, processing vast amounts of data in real-time requires high-performance computing environments that can handle the intensive workloads generated by these sophisticated models. This requirement may lead to significant energy consumption and increased operational costs, which could be prohibitive.
Another critical limitation is the adaptability of these systems to different network environments. The testbed configurations, which often include a variety of network scenarios and attack vectors, must be meticulously designed to mirror the diverse real-world network landscapes. However, this can be challenging as each network environment has unique characteristics and threat profiles that may not be fully addressed by a standardized IDS solution. Moreover, maintaining the accuracy of such systems in dynamic environments where new threats continuously emerge is problematic. The models must be constantly updated and trained on the latest threat data to remain effective, which involves continuous data collection and analysis, further straining resources.
@@ 7.2 LLMs challenges
(f) Transfer learning from pre-trained LLM: Pre-trained models such as GPT and BERT reveal a range of limitations and potential biases that are critical to consider. One of the primary limitations is the heavy reliance on extensive and diverse datasets for training these models. While pre-trained models have demonstrated remarkable capabilities in understanding and generating text based on their training, they are often constrained by the data they were trained on. This dependency can lead to biases in model outputs, particularly if the training data was not balanced or if it contained inherent prejudices. For instance, models like GPT and BERT may inadvertently perpetuate or amplify biases present in their training sets, leading to skewed or unfair outcomes when deployed in real-world applications. Moreover, these models require significant computational resources not only for training but also for fine-tuning and deployment, which can be a barrier for institutions with limited hardware capabilities. This creates a divide in who can effectively use these technologies, potentially widening the gap between large organizations and smaller entities.
Another concern is the adaptability of these models to specific tasks or domains. While they are generally powerful, transferring a pre-trained model to a specialized domain without extensive retraining can lead to suboptimal performance. This limitation necessitates additional layers of fine-tuning, which again demands more data and computational power, further complicating their use in varied applications. These challenges highlight the need for ongoing research to mitigate biases, reduce resource demands, and enhance the flexibility of pre-trained models for broader and more equitable usage.
@@ 7.3 Perspectives and future directions
Since LLMs represent the latest advancement in Transformer technology, future research and perspectives mostly focus exclusively on LLMs, but are also applicable to Transformers. Despite extensive research into LLMs within cybersecurity, their exploration and application are still in the early stages, offering considerable potential for growth [155]. The complexity of cybersecurity arises not only from the variety of attack methods but also from the intricate nature of network environments, combined with the necessity for a comprehensive application of diverse tools and strategies to ensure effective protection [156, 21]. Addressing these challenges requires AI systems with advanced capabilities in planning, reasoning, tool use, memory, and more. Building upon this discussion, the authors identify several critical perspectives that need to be addressed in LLM-based IDS, as follows:
@@
(g) Integration with dynamic security Ecosystems and adversarial resilience: Future LLMs for IDS need to increasingly integrate with dynamic security ecosystems, incorporating real-time threat intelligence from diverse sources such as threat feeds, vulnerability databases, and incident reports. For example, the suggested EvolveDroid system [174] enhances LLM ecosystem security by leveraging multiple feature types and contrastive learning to enhance detection capabilities and adapt to evolving threats. This integration not only enriches model training but also enhances adversarial resilience by enabling LLMs to detect and mitigate sophisticated adversarial attacks. Additionally, collaborative defense systems where multiple specialized LLMs work synergistically will further strengthen IDS capabilities, improving overall threat detection accuracy and reducing false positives. This holistic approach ensures that LLM-based IDS are well-equipped to handle evolving cybersecurity challenges effectively. Nevertheless, the field is still in the early stages of research investigation and requires further exploration and development to fully understand its implications and potential.
(h) Cross-domain applications of Transformer and LLM-based IDS: As highlighted by the limited cross-domain efficacy of traditional DL-based NIDS models [175, 176], there is a need for more resilient and flexible architectures. The future direction of utilizing Transformers and LLMs in IDS shows promise for cross-domain applications, especially in sectors like healthcare and finance, which face significant network security challenges [177]. In healthcare, fine-tuned LLMs [178] can be specifically tailored to detect and analyze anomalies and potential threats in sensitive medical data, ensuring patient data integrity and confidentiality. This is critical as the sector handles vast amounts of private information that, if compromised, could have severe consequences. Utilizing Task-Fine-Tune allows these models to precisely identify and react to unique patterns of cyber threats that are specific to the healthcare domain, such as ransomware attacks on hospital systems or unauthorized access to medical records. Similarly in finance [177], LLMs optimized for IDS tasks can monitor and analyze transactional data for signs of fraudulent activities with heightened accuracy. Financial institutions are often targets of sophisticated cyber-attacks aimed at financial gain. By fine-tuning LLMs on specific financial datasets, such as those involving transactional behaviors and fraud indicators, these models can become highly effective at detecting and mitigating potential threats, thereby safeguarding assets and sensitive financial information. Both sectors benefit from the ability of LLMs and Transformers to process and analyze large volumes of text and other data types quickly and accurately. This capability allows for real-time threat detection and enhances the overall security infrastructure of organizations in these critical fields. By incorporating these advanced technologies, healthcare and finance institutions can better protect sensitive data and ensure compliance with regulatory standards, ultimately safeguarding both individual privacy and financial integrity. However, the use of LLM-based IDS in healthcare and finance remains an under-researched area that warrants further exploration in future studies.
@@
(i) Interpretability, scalability, and real-time improvement: Integrating domain knowledge into Transformer and LLM-based IDS offers a promising direction for future research. By incorporating specific insights and contextual information from various fields, the interpretability of these models can be significantly improved. This approach allows for a deeper understanding of the patterns and anomalies detected, making the results more meaningful and actionable for cybersecurity experts. For instance, in the healthcare sector, integrating medical domain knowledge can help the model distinguish between normal and abnormal network activities related to medical devices and patient data flows. Similarly, in the financial sector, understanding the typical transaction patterns and identifying deviations can enhance the model’s ability to detect fraudulent activities.
@@
Moreover, the development of pre-trained Transformers and LLMs tailored for IDS facilitates the creation of lightweight models adept at real-time threat processing. For example, small language models such as TinyLlama [181] offer compact versions of LLMs, engineered to provide similar capabilities for specialized tasks with significantly reduced computational and memory demands. These models proficiently process and interpret both text and tabular data. In real-time intrusion processing, SLMs could be used to rapidly analyze vast quantities of log data, enabling quicker and more effective threat detection with minimal hardware investment. This expedited processing capability is crucial for immediate threat responses in cybersecurity environments. Researchers are urged to explore the potential of real-time Transformers and LLMs-based IDS by applying their methods on embedded platforms such as field programmable gate arrays, and by measuring real-time metrics including inference time, latency, resource utilization, and algorithmic complexity. Such evaluations provide valuable insights for users and developers, helping determine the most suitable Transformer or LLM for specific hardware setups and IDS threat types.